Moment of inertia of a uniform circular disc about its diameter is l. Its moment of inertia about an axis parallel to its plane and passing through a point on its rim will be what???

Grade:11

2 Answers

Aakash
6 years ago
About diameter =(MR^2)/4And about point passing through rim:Using parallel axis theorem:(MR^2)/4 + MR^2=(5MR^2)/4
chandana chandhu
5 years ago
By the principle of parallel axis,
I={MR^2/4}+MR^2=5MR^2/4,for a tangent parallel to diameter
For a tangent parallel to axis passing through its center and perpendicular to its plane 
I={MR^2/2}+MR^2=3MR^2/2